Six years hence a man?s age will be three times his son!s age and three years ago he was nine times as old as his son . Find their present ages?
Answer by JulietG(1812)

M+6=3(S+6)
Distribute
M+6=3S+18
Subtract 6 from each side
M = 3S+12


M-3 = 9(S-3)
Distribute
M-3=9s-27
Add 3 to each side
M = 9S-24


Take the value of M from the first equation and substitute it for M in the second.
M = 9S-24
(3S+12)=9S-24
Subtract 3S from each side
12 = 6S-24
Add 24 to each side
36 = 6S
Divide each side by 6
S = 6


If the son is 6 years old, then 3 years ago he was 3 years old. At that time, his dad was 9x his age, or 27. That was 3 years ago. Now he's 30.
Son=6
Dad=30